OpenStudy (anonymous):

laplace transform of sin^3 2t

OpenStudy (anonymous):

Here are some basic transforms you can use, http://image.mathcaptain.com/cms/images/83/laplace-tranform-formulas-list.png

OpenStudy (anonymous):

thanks ,tried using them,but the main problem is coming from the power rised(3)

zepdrix (zepdrix):

\[\large \mathscr{L}\left[\sin^3(2t)\right] \qquad = \qquad \mathscr{L}\left[\sin(2t)\sin^2(2t)\right] \qquad =\qquad \\ \large \mathscr{L}\left[\sin(2t)\left(1-\cos^2(2t)\right)\right] \qquad = \qquad \mathscr{L}\left[\sin(2t)\right]-\mathscr{L}\left[\sin(2t)\cos^2(2t)\right]\]

zepdrix (zepdrix):

Woops it got cut off..\[\large \mathscr{L}\left[\sin(2t)\right]-\mathscr{L}\left[\sin(2t)\cos^2(2t)\right]\] That first term isn't too bad right? For the other one, use the definition of the Laplace Transform. You'll have to do integration by parts a couple times.

OpenStudy (amistre64):

\[L\{sin^3(2t)\}=\int_{0}^{inf}e^{-st}sin^3(2t)~dt\] which still amounts to some by parts stuff

OpenStudy (anonymous):

I propose doing: \[\mathcal{L} \left( \sin^3(2t) \right)=\int\limits_0^{\infty}e^{-st} \left( \frac{e^{2ti}-e^{-2ti}}{2i}\right)^3dt\] This **should** make your life easier.